Brand X sells 21 oz bags of mixed nuts that contain 29% peanuts. To make their product they combine Brand A mixed nuts which contain 35% peanuts and Brand B mixed nuts which contain 25% peanuts. How much of each do they need to use?
Answer by ptaylor(2198)

Let x=Amount of Brand A
Then 21-x=Amount of Brand B
Amount of peanuts in Brand A and B before the mixture takes place has to equal the amount of peanuts in the final mixture, so our equation to solve is:
0.35x+0.25(21-x)=21*0.29 get rid of parens (distributive law)
0.35x+5.25-0.25x=6.09 subtract 5.25 from each side
0.35x+5.25-5.25-0.25x=6.09-5.25 collect like terms
0.10x=0.84 divide each side by 0.10
x=8.4 oz---------------------------------Amount of Brand A needed
21-x=21-8.4=12.6-----------------------------Amount of Brand B needed
CK
8.4*0.35+12.6*0.25=21*0.29
2.94+3.15=6.09
6.09=6.09
Hope this helps---ptaylor